# Spin-dependent Force from an NV center Ensemble on a Microlever

### v1

@@ E.1 Dynamical equations of the spin
We restrict the calculation to the two level subsystem composed of the |$\ket{0^{\prime}}$ state and either the |$\ket{-1^{\prime}}$ or the |$\ket{+1^{\prime}}$ state. For simplicity of notations, we write |$\ket{g}$ the ground state of this two level-system, and |$\ket{e}$ the excited state. By considering the density matrix formalism, we can derive the master equation:
@@
where $\rho_{\rm ee}$ (resp. $\rho_{\rm gg}$) is the population in state |$\ket{e}$ (resp. |$\ket{g}$) verifying $\rho_{\rm ee}+\rho_{\rm gg}\approx 1$, $\rho_{\rm eg}$ are the coherences between these two states in the frame rotating at the energy transitions of the two level-system. $\Omega$ is the Rabi frequency. We neglect the influence of the mechanical displacement in the detuning $\Delta(t)$, since the displacement of the mechanical oscillator is too small to observe back-action from the spin towards the dynamics as was observed in Delord et al. (2020).
@@ E.2 Dynamic equation of the cantilever mechanical mode
where $F_{y}=\frac{\tau_{e,y}-\tau_{g,y}}{l_{\rm e}}$, $\tau_{e,y}$ (resp. $\tau_{g,y}$) is the torque assuming $100\%$ population in the |$\ket{e}$ state (resp. |$\ket{g}$ state).
E.3 Calculation of the quadrature at the mechanical resonance
@@
Appendix F FM-MDMR with substraction of signals obtained with and without green laser
@@
In Fig. 8-(a), we present two FM-MDMR scans taken on the Y quadrature at a microwave power of $30~{}{\rm dBm}$. The red curve shows the cantilever displacement in the presence of a green laser excitation, while the blue curve was obtained without green laser. The magnetic resonance of the spin transition is equal to $2.962~{}{\rm GHz}$. The FM-MDMR of Fig. 8-(b), shows the subtraction of the two curves. We observe an expected dispersive curve centered on the magnetic resonance frequency.
@@
Figure 8: (a) FM-MDMR measured on the Y quadrature as a function of the microwave frequency under green laser excitation (red curve) and without green laser excitation (blue curve) (b) Subtracted FM-MDMR as a function of the microwave frequency.

### v2

@@ E.1 Dynamical equations of the spin
We restrict the calculation to the two-level subsystem composed of the $|{0^{\prime}}\rangle$ state and either the $|{-1^{\prime}}\rangle$ or the $|{+1^{\prime}}\rangle$ state. For simplicity of notations, we write $|{g}\rangle$ the ground state of this two level-system, and $|{e}\rangle$ the excited state. By considering the density matrix formalism, we can derive the master equation:
@@
where $\rho_{\rm ee}$ (resp. $\rho_{\rm gg}$) is the population in state $|{e}\rangle$ (resp. $|{g}\rangle$) verifying $\rho_{\rm ee}+\rho_{\rm gg}\approx 1$, $\rho_{\rm eg}$ are the coherences between these two states in the frame rotating at the energy transitions of the two level-system. $\Omega$ is the Rabi frequency. We neglect the influence of the mechanical displacement in the detuning $\Delta(t)$, since the displacement of the mechanical oscillator is too small to observe back-action from the spin towards the dynamics as was observed in Delord et al. (2020).
@@ E.2 Dynamic equation of the cantilever mechanical mode
where $F_{y}=\frac{\tau_{e,y}-\tau_{g,y}}{l_{\rm e}}$, $\tau_{e,y}$ (resp. $\tau_{g,y}$) is the torque assuming $100\%$ population in the $|{e}\rangle$ state (resp. $|{g}\rangle$ state).
E.3 Quadratures’ amplitudes at the mechanical resonance
@@
Annexe F FM-MDMR with subtraction of signals obtained with and without green laser
@@
In Fig. 8-(a), we present two FM-MDMR scans taken on the $Y$ quadrature at a microwave power of $30\leavevmode\nobreak\ {\rm dBm}$. The red curve shows the cantilever displacement in the presence of a green laser excitation, while the blue curve was obtained without green laser. The magnetic resonance of the spin transition is equal to $2.962\leavevmode\nobreak\ {\rm GHz}$. The FM-MDMR of Fig. 8-(b), shows the subtraction of the two curves. We observe an expected dispersive curve centered on the magnetic resonance frequency.
@@
Figure 8: (a) FM-MDMR measured on the $Y$ quadrature as a function of the microwave frequency under green laser excitation (red curve) and without green laser excitation (blue curve). (b) Subtracted FM-MDMR as a function of the microwave frequency.
Références
  • Aspelmeyer et al. (2014) Markus Aspelmeyer, Tobias J Kippenberg,  and Florian Marquardt, “Cavity optomechanics,” Reviews of Modern Physics 86, 1391–1452 (2014).
  • Barzanjeh et al. (2022) Shabir Barzanjeh, André Xuereb, Simon Gröblacher, Mauro Paternostro, Cindy A Regal,  and Eva M Weig, “Optomechanics for quantum technologies,” Nature Physics 18, 15–24 (2022).
  • Mercier de Lépinay et al. (2021) Laure Mercier de Lépinay, Caspar F Ockeloen-Korppi, Matthew J Woolley,  and Mika A Sillanpää, “Quantum mechanics–free subsystem with mechanical oscillators,” Science 372, 625–629 (2021).
  • Kotler et al. (2021) Shlomi Kotler, Gabriel A Peterson, Ezad Shojaee, Florent Lecocq, Katarina Cicak, Alex Kwiatkowski, Shawn Geller, Scott Glancy, Emanuel Knill, Raymond W Simmonds, et al., “Direct observation of deterministic macroscopic entanglement,” Science 372, 622–625 (2021).
  • Bild et al. (2023) Marius Bild, Matteo Fadel, Yu Yang, Uwe Von Lüpke, Phillip Martin, Alessandro Bruno,  and Yiwen Chu, “Schrödinger cat states of a 16-microgram mechanical oscillator,” Science 380, 274–278 (2023).
  • Rabl et al. (2009) Peter Rabl, P Cappellaro, MV Gurudev Dutt, Liang Jiang, JR Maze,  and Mikhail D Lukin, “Strong magnetic coupling between an electronic spin qubit and a mechanical resonator,” Physical Review B 79, 041302 (2009).
  • Rugar et al. (2004) Daniel Rugar, Raffi Budakian, HJ Mamin,  and BW Chui, “Single spin detection by magnetic resonance force microscopy,” Nature 430, 329–332 (2004).
  • Rabl (2010) P Rabl, “Cooling of mechanical motion with a two-level system: The high-temperature regime,” Physical Review B—Condensed Matter and Materials Physics 82, 165320 (2010).
  • Rusconi et al. (2022) Cosimo C Rusconi, Maxime Perdriat, Gabriel Hétet, Oriol Romero-Isart,  and Benjamin A Stickler, “Spin-controlled quantum interference of levitated nanorotors,” Physical Review Letters 129, 093605 (2022).
  • Yin et al. (2013) Zhang-qi Yin, Tongcang Li, Xiang Zhang,  and LM Duan, “Large quantum superpositions of a levitated nanodiamond through spin-optomechanical coupling,” Physical Review A 88, 033614 (2013).
  • Ma et al. (2017) Yue Ma, Thai M Hoang, Ming Gong, Tongcang Li,  and Zhang-qi Yin, “Proposal for quantum many-body simulation and torsional matter-wave interferometry with a levitated nanodiamond,” Physical Review A 96, 023827 (2017).
  • Wei et al. (2015) Bo-Bo Wei, Christian Burk, Jörg Wrachtrup,  and Ren-Bao Liu, “Magnetic ordering of nitrogen-vacancy centers in diamond via resonator-mediated coupling,” EPJ Quantum Technology 2, 1–7 (2015).
  • Delord et al. (2017) T Delord, L Nicolas, Y Chassagneux,  and G Hétet, “Strong coupling between a single nitrogen-vacancy spin and the rotational mode of diamonds levitating in an ion trap,” Physical Review A 96, 063810 (2017).
  • Ovartchaiyapong et al. (2014) Preeti Ovartchaiyapong, Kenneth W. Lee, Bryan A. Myers,  and Ania C. Bleszynski Jayich, “Dynamic strain-mediated coupling of a single diamond spin to a mechanical resonator,” Nature Communications 5, 4429 (2014).
  • Teissier et al. (2014) J. Teissier, A. Barfuss, P. Appel, E. Neu,  and P. Maletinsky, “Strain coupling of a nitrogen-vacancy center spin to a diamond mechanical oscillator,” Phys. Rev. Lett. 113, 020503 (2014).
  • Arcizet et al. (2011) Olivier Arcizet, Vincent Jacques, Alessandro Siria, Philippe Poncharal, Pascal Vincent,  and Signe Seidelin, “A single nitrogen-vacancy defect coupled to a nanomechanical oscillator,” Nature Physics 7, 879–883 (2011).
  • Kolkowitz et al. (2012) Shimon Kolkowitz, Ania C Bleszynski Jayich, Quirin P Unterreithmeier, Steven D Bennett, Peter Rabl, JGE Harris,  and Mikhail D Lukin, “Coherent sensing of a mechanical resonator with a single-spin qubit,” Science 335, 1603–1606 (2012).
  • Delord et al. (2020) T. Delord, P. Huillery, L. Nicolas,  and G. Hétet, “Spin-cooling of the motion of a trapped diamond,” Nature 580, 56–59 (2020).
  • Perdriat et al. (2022) Maxime Perdriat, Paul Huillery, Clément Pellet-Mary,  and Gabriel Hétet, “Angle locking of a levitating diamond using spin diamagnetism,” Physical Review Letters 128, 117203 (2022).
  • Rabe et al. (1996) U Rabe, K Janser,  and Wt Arnold, “Vibrations of free and surface-coupled atomic force microscope cantilevers: Theory and experiment,” Review of scientific instruments 67, 3281–3293 (1996).
  • Fischer et al. (2019) Ran Fischer, Dylan P McNally, Chris Reetz, Gabriel GT Assumpcao, T Knief, Yiheng Lin,  and Cindy A Regal, “Spin detection with a micromechanical trampoline: towards magnetic resonance microscopy harnessing cavity optomechanics,” New Journal of Physics 21, 043049 (2019).
  • Fogliano et al. (2021) Francesco Fogliano, Benjamin Besga, Antoine Reigue, Laure Mercier de Lépinay, Philip Heringlake, Clement Gouriou, Eric Eyraud, Wolfgang Wernsdorfer, Benjamin Pigeau,  and Olivier Arcizet, “Ultrasensitive nano-optomechanical force sensor operated at dilution temperatures,” Nature Communications 12, 4124 (2021).
  • Doherty et al. (2013) Marcus W. Doherty, Neil B. Manson, Paul Delaney, Fedor Jelezko, Jörg Wrachtrup,  and Lloyd C.L. Hollenberg, “The nitrogen-vacancy colour centre in diamond,” Physics Reports 528, 1–45 (2013).
  • Cleland (2013) Andrew N Cleland, Foundations of nanomechanics: from solid-state theory to device applications (Springer Science & Business Media, 2013).
  • Mehlin et al. (2015) A Mehlin, F Xue, D Liang, HF Du, MJ Stolt, S Jin, ML Tian,  and M Poggio, “Stabilized skyrmion phase detected in mnsi nanowires by dynamic cantilever magnetometry,” Nano letters 15, 4839–4844 (2015).
  • Gross et al. (2016) B Gross, DP Weber, D Rüffer, A Buchter, F Heimbach, A Fontcuberta i Morral, D Grundler,  and M Poggio, “Dynamic cantilever magnetometry of individual cofeb nanotubes,” Physical Review B 93, 064409 (2016).
  • Gross et al. (2020) Boris Gross, Simon Philipp, Korbinian Geirhos, Andrea Mehlin, Sándor Bordács, Vladimir Tsurkan, A Leonov, István Kézsmárki,  and Martino Poggio, “Stability of néel-type skyrmion lattice against oblique magnetic fields in GaV4S8 and GaVSe8,” Physical Review B 102, 104407 (2020).
  • Gross et al. (2021) B Gross, S Philipp, E Josten, Jonathan Leliaert, Erik Wetterskog, Lennart Bergström,  and M Poggio, “Magnetic anisotropy of individual maghemite mesocrystals,” Physical Review B 103, 014402 (2021).
  • Kaviani et al. (2023) Hamidreza Kaviani, Bishnupada Behera, Ghazal Hajisalem, Gustavo de Oliveira Luiz, David P. Lake,  and Paul E. Barclay, “High-frequency torsional motion transduction using optomechanical coupled oscillators,” Optica 10, 35–41 (2023).
  • Kim et al. (2016) P. H. Kim, B. D. Hauer, C. Doolin, F. Souris,  and J. P. Davis, “Approaching the standard quantum limit of mechanical torque sensing,” Nature Communications 7, 13165 (2016).
  • Einstein and De Haas (1915) A Einstein and WJ De Haas, “Experimental proof of the existence of ampère’s molecular currents,” in Proc. KNAW, Vol. 181 (1915) p. 696.
  • Mori et al. (2020) K Mori, MG Dunsmore, JE Losby, DM Jenson, M Belov,  and MR Freeman, “Einstein–de haas effect at radio frequencies in and near magnetic equilibrium,” Physical Review B 102, 054415 (2020).
  • Alzetta et al. (1967) G Alzetta, Ennio Arimondo, C Ascoli,  and A Gozzini, “Paramagnetic resonance experiments at low fields with angular-momentum detection,” Il Nuovo Cimento B (1965-1970) 52, 392–402 (1967).
  • Arimondo (1968) Ennio Arimondo, “Angular momentum detection of non-linear phenomena in paramagnetic resonance,” in Annales de Physique, Vol. 14 (EDP Sciences, 1968) pp. 425–447.
  • Pellet-Mary et al. (2021) C Pellet-Mary, Paul Huillery, M Perdriat,  and G Hétet, “Magnetic torque enhanced by tunable dipolar interactions,” Physical Review B 104, L100411 (2021).
  • Perdriat (2023) Maxime Perdriat, Angular control of levitated magnetic particles for studies of gyromagnetism at the micro-scale, Ph.D. thesis, Université paris Cité (2023).
  • Perdriat et al. (2021) Maxime Perdriat, Clément Pellet-Mary, Paul Huillery, Loïc Rondin,  and Gabriel Hétet, “Spin-mechanics with nitrogen-vacancy centers and trapped particles,” Micromachines 12 (2021), 10.3390/mi12060651.
  • Losby et al. (2018) Joseph E Losby, Vincent TK Sauer,  and Mark R Freeman, “Recent advances in mechanical torque studies of small-scale magnetism,” Journal of Physics D: Applied Physics 51, 483001 (2018).